OpenStudy (anonymous):

A snail moves 4 metres along a straight line in 1 hour and in each succeeding hour 20% less than the preceeding one.find how far it goes.

OpenStudy (anonymous):

no time limit????

OpenStudy (anonymous):

20 metres

OpenStudy (anonymous):

GP a= 4, r= 0.8 S=a/ (1-r) = 4/0.2

OpenStudy (anonymous):

= 4 x 5 =20

OpenStudy (anonymous):

thanks